# Image Coding for Machines with Edge Information Learning Using Segment Anything

### v1

@@ Abstract
Image Coding for Machines (ICM) is an image compression technique for image recognition. This technique is essential due to the growing demand for image recognition AI. In this paper, we propose a method for ICM that focuses on encoding and decoding only the edge information of object parts in an image, which we call SA-ICM. This is an Learned Image Compression (LIC) model trained using edge information created by Segment Anything. Our method can be used for image recognition models with various tasks. SA-ICM is also robust to changes in input data, making it effective for a variety of use cases. Additionally, our method provides benefits from a privacy point of view, as it removes human facial information on the encoder’s side, thus protecting one’s privacy. Furthermore, this LIC model training method can be used to train Neural Representations for Videos (NeRV), which is a video compression model. By training NeRV using edge information created by Segment Anything, it is possible to create a NeRV that is effective for image recognition (SA-NeRV). Experimental results confirm the advantages of SA-ICM, presenting the best performance in image compression for image recognition. We also show that SA-NeRV is superior to ordinary NeRV in video compression for machines.
@@ 1 Introduction
In this paper, we propose a new ICM model (SA-ICM) that solves the problems of the above approaches. The proposed method is a type of RL-based approach, which does not require additional information such as ROI-map as encoder input and does not learn using task-loss. On the other hand, unlike existing RL-based methods, it does not use mask images in the COCO dataset. Instead, SA-ICM uses mask images generated by Segment Anything [17]. The edges of the segmentation map generated using Segment Anything are used to train the LIC model. This creates an LIC model that can decode only the main edge information. The proposed method reduces more textures than existing RL-based approaches and while also removing human face textures. It has good properties both in terms of image compression and privacy protection. Furthermore, this learning method can be applied to the NeRV [18] learning method to create a video compression model for image recognition (SA-NeRV). In experiments, we will investigate the image compression performance of SA-ICM for image recognition and compare it to other methods for ICM to demonstrate the effectiveness of this method. For the image recognition models, we use an object detection model, an instance segmentation model, and a panoptic segmentation model to show the robustness to changes in the image recognition model. By using COCO, VisDrone [27], and Cityscapes [20] as the datasets, we show that our method can be used in various use cases. We also compare the image recognition accuracy between SA-NeRV and ordinary NeRV decoded images to confirm the effectiveness of SA-NeRV.
@@ 3.1 SA-ICM
We create a method to encode and decode only the edge information in an image by training the LIC model using a segment anything model (SAM). This method is a variant of the RL-based approach, which requires no additional information input and does not train the model using task-loss. The original RL-based approach [15] used the image shown in Fig.2(b) as the mask image and trained the LIC model only on the object regions in the image. However, the decoded image using this approach is not suitable for image recognition tasks that require background information in the image because the background in the image is represented roughly. In addition, unnecessary textures of object parts are decoded, meaning there are still rooms for improvement in compression performance.
@@ 3.2 SA-NeRV
In (5), $x$, $\hat{x}$, $T$ and $\beta$ have the same meaning as those variables and constants in (4). $\mathcal{L}_{n}$ is the loss component in NeRV training, as shown in Eq. (4). This method can be used to decode images that are useful for image recognition models.
@@ 4.3 Experimental Methods for Evaluating SA-NeRV
To evaluate the efficiency of SA-NeRV, we compare the image recognition accuracy of the original NeRV decode images and SA-NeRV decoded images. We use the SFU-HW-Objects-v1 [27] dataset, consisting 18 sequences and their corresponding annotations for object detection. These sequences are classified into five classes (A to E) according to the image size and the characteristics. This dataset is also applied for the Common Test Condition in MPEG’s VCM standardization activities. In this experiment, we work with class C and class D sequences from this data set. After embedding these videos in NeRV and SA-NeRV, the videos are decoded. We use the pre-trained YOLOv7 to measure the object detection accuracy in the decoded videos.
@@ 4.4 SA-NeRV Evaluation Experimental Results
  sequence name
 NeRV [18]
 SA-NeRV
 BQMall
  28.03
  28.24
 BasketballDrill
  34.26
  34.93
 PartyScene
  34.34
  34.61
 RaceHorsesC
  80.99
  81.77
 BQSquare
  27.84
  29.80
 BasketballPass
  23.29
  24.88
 BlowingBubbles
  41.83
  48.84
 RaceHorsesD
  89.12
  88.98

### v3

@@ Abstract
Image Coding for Machines (ICM) is an image compression technique for image recognition. This technique is essential due to the growing demand for image recognition AI. In this paper, we propose a method for ICM that focuses on encoding and decoding only the edge information of object parts in an image, which we call SA-ICM. This is an Learned Image Compression (LIC) model trained using edge information created by Segment Anything. Our method can be used for image recognition models with various tasks. SA-ICM is also robust to changes in input data, making it effective for a variety of use cases. Additionally, our method provides benefits from a privacy point of view, as it removes human facial information on the encoder’s side, thus protecting one’s privacy. Furthermore, this LIC model training method can be used to train Neural Representations for Videos (NeRV), which is a video compression model. By training NeRV using edge information created by Segment Anything, it is possible to create a NeRV that is effective for image recognition (SA-NeRV). Experimental results confirm the advantages of SA-ICM, presenting the best performance in image compression for image recognition. We also show that SA-NeRV is superior to ordinary NeRV in video compression for machines. Code is available at https://github.com/final-0/SA-ICM.
@@ 1 Introduction
In this paper, we propose a new ICM model (SA-ICM) that solves the problems of the above approaches. The proposed method is a type of RL-based approach, which does not require additional information such as ROI-map as encoder input and does not learn using task-loss. On the other hand, unlike existing RL-based methods, it does not use mask images in the COCO dataset. Instead, SA-ICM uses mask images generated by Segment Anything Model (SAM) [17]. The edges of the segmentation map generated using SAM are used to train the LIC model. This creates an LIC model that can decode only the main edge information. The proposed method reduces more textures than existing RL-based approaches and while also removing human face textures. It has good properties both in terms of image compression and privacy protection. Furthermore, this learning method can be applied to the NeRV [18] learning method to create a video compression model for image recognition (SA-NeRV). In experiments, we will investigate the image compression performance of SA-ICM for image recognition and compare it to other methods for ICM to demonstrate the effectiveness of this method. For the image recognition models, we use an object detection model, an instance segmentation model, and a panoptic segmentation model to show the robustness to changes in the image recognition model. By using COCO, VisDrone [27], and Cityscapes [20] as the datasets, we show that our method can be used in various use cases. We also compare the image recognition accuracy between SA-NeRV and ordinary NeRV decoded images to confirm the effectiveness of SA-NeRV.
@@ 3.1 SA-ICM
We propose SA-ICM, a method to encode and decode only the edge information in an image by training the LIC model using a segment anything. This method is a variant of the RL-based approach, which requires no additional information input and does not train the model using task-loss. The original RL-based approach [15] used the image shown in Fig.2(b) as the mask image and trained the LIC model only on the object regions in the image. However, the decoded image using this approach is not suitable for image recognition tasks that require background information in the image because the background in the image is represented roughly. In addition, unnecessary textures of object parts are decoded, meaning there are still rooms for improvement in compression performance.
@@ 3.2 SA-NeRV
In (6), $x$, $\hat{x}$, $T$ and $\beta$ have the same meaning as those variables and constants in (4). $\mathcal{L}_{n}$ is the loss component in NeRV training, as shown in Eq. (4). By learning NeRV with $\mathcal{L}_{sa-n}$ as the loss function, the position and shape of objects in the video are efficiently embedded in the neural network. This method can be used to decode images that are useful for image recognition models.
@@ 4.3 Experimental Methods for Evaluating SA-NeRV
To evaluate the efficiency of SA-NeRV, we compare the image recognition accuracy of the original NeRV decode images and SA-NeRV decoded images. We use the SFU-HW-Objects-v1 [27] dataset, consisting 18 sequences and their corresponding annotations for object detection. These sequences are classified into five classes (A to E) according to the image size and the characteristics. This dataset is also applied for the Common Test Condition in MPEG’s VCM standardization activities. In this experiment, we work with class C and class D sequences from this data set. After embedding these videos in NeRV and SA-NeRV, the videos are decoded. We use the pre-trained YOLOv7 [28] to measure the object detection accuracy in the decoded videos.
@@ 4.4 SA-NeRV Evaluation Experimental Results
  sequence name  NeRV [18]  SA-NeRV
 BQMall   28.03   28.24
 BasketballDrill   34.26   34.93
 PartyScene   34.34   34.61
 RaceHorsesC   80.99   81.77
 BQSquare   27.84   29.80
 BasketballPass   23.29   24.88
 BlowingBubbles   41.83   48.84
 RaceHorsesD   89.12   88.98